# Chemistry Beyond Exact Solutions on a Quantum-Centric Supercomputer

### v2

@@ V.3 Effect of orbital optimizations in the dissociation of N2 (cc-pVDZ)
Figure S15: Effect of orbital optimizations (OO) in the performance of the HPC quantum estimator with configuration recovery in the dissociation of N2 (cc-pVDZ). The ground-state energy is shown as a function of the bondlength. The estimator is run on measurement outcomes from a quantum device. The energies shown correspond to the lowest energy amongst the $K$ batches of configurations, i.e. ${\min_{k}\left(E^{(k)}\right)}$. Energies from different classical methods are shown for reference, as indicated in the legend. Results are shown for the estimator run with different numbers of configurations $d$, as indicated in each panel (a)-(d).
Figure S16: Orbital occupancy profile ${\bf{n}}$ for the [2Fe-2S], produced by the HPC quantum estimator with configuration recovery. The increasing index $p$ labels molecular orbitals of increasing energy. The occupancy profile is shown for the three eigenstates identified in the energy-variance analysis (see Fig. 4 (b) in the main text).
Figure S17: Orbital occupancy profile $\tilde{{\bf{n}}}$ for the [2Fe-2S], produced by the HPC quantum estimator with configuration recovery. The increasing index $t$ labels localized orbitals as shown in the bottom of the panel. The occupancy profile is shown for the three eigenstates identified in the energy-variance analysis (see Fig. 4 (b) in the main text). The representation of the orbitals is obtained from the repository [108] from Ref. [45].
V.4 Orbital occupation numbers for the low-energy spectrum of [2Fe-2S]
The calculations reported in Figure 4 of the main text, for the [2Fe-2S] cluster, identified three distinct eigenstates that we labeled A, B, and C. In this section, we provide additional information about the nature of the three identified eigenstates. The purpose of this analysis is to illustrate the robust classification of the wavefunctions produced by our estimator into three families (A, B, and C) but also to shed light on the nature of the corresponding eigenstates and the ability of our method to accurately capture a multireference character in electronic eigenstates, pinpointing areas of possible future improvement. In particular, we show the spatial orbital occupancy for the three eigenstates that the HPC quantum estimator finds for [2Fe-2S] both in the MO basis and in a basis of spatially local orbitals. The MO basis helps highlight deviations from the mean-field states or lack thereof, and the basis of localized orbitals (identifiable with e.g. Fe $3d$ and S $2p$ orbitals) helps characterize our method’s ability to capture local antiferromagnetic correlations between transition-metal atoms.
@@ V.4 Orbital occupation numbers for the low-energy spectrum of [2Fe-2S]
To further understand the nature of Eigenstates B and C, we study the occupancy in the basis of localized orbitals $\tilde{{\bf{n}}}$ (see Sec. I.2 for details about the basis of localized orbitals). The reported occupancies are constructed as follows: first, we obtain the one-body density matrix for the approximate ground state obtained from each batch of configurations, expressed in the basis of MOs: $\Gamma_{pq,\sigma}^{(k)}$ (see Eq. S41). We then rotate the density matrix to the basis of localized orbitals with a similarity transformation denoted by $\Omega$,
@@
Fig. S17 shows a comparison of the value of $\tilde{{\bf{n}}}$ for Eigenstates A, B, and C. The reflection symmetry along the plane that separates the left and right sides of the molecule is broken for the three approximate eigenstates. This is highlighted by the observation that the occupancy of the equivalent orbitals for the two Fe atoms are not the same. We also observe that the occupancy of the bonding orbitals of the SCH3 groups and the occupancy of the $3p$ orbitals of the S atoms is similar for the three eigenstates. The occupancy of the $3d$ orbitals of Fe has the largest differences among the three different eigenstates. The more notable discrepancy in the d-orbitals of Fe may be a consequence of the fact that the low-energy physics of this molecule is dominated by the antiferromagnetic coupling of the electrons in the Fe $3d$ orbitals [45, 4].
V.5 Quantum signal in the experiments
@@
one should verify whether there is a useful signal coming out of the quantum circuits, comparing samples from $\tilde{P}_{\Psi}$ and uniform random samples. We first investigate the fraction of sampled configurations that live in the correct particle sector $p_{s}^{\textrm{hw}}$ and compare it to the fraction that would be obtained if the samples came from the uniform distribution $p_{s}^{\textrm{unif}}$. The values for $p_{s}^{\textrm{hw}}$ alongside their 95% confidence interval, and $p_{s}^{\textrm{unif}}$ are reported in Table 2. For the 40-, 52-, and 72-qubit experiments, we observe that $p_{s}^{\textrm{hw}}$ is $\sim 20$, $\sim 10^{3}$, and $\sim 40$ times larger than $p_{s}^{\textrm{unif}}$ respectively. This provides a first indication that configurations sampled from the quantum device contain a signal that is distinguishable from white noise.
The second set of tests that we conduct is the comparison of the ground-state energy obtained by the HPC quantum estimator with and without applying the configuration recovery procedure. Note that not applying configuration recovery makes the estimator equivalent to the QSCI framework [30] . Having reasonable energies (lower than or in the vicinity of the RHF energy) without applying configuration recovery reveals that the set $\mathcal{X}_{N}$ already contains reasonable configurations to construct a non-random ground state.
The third set of tests involves comparing the performance of the HPC quantum estimator with configuration recovery run on configurations obtained from the quantum device and samples obtained from the uniform distribution. We consider the uniform distribution over all possible configurations in the Fock space, as well as the uniform distribution over configurations with the correct particle number, as in Fig. 4 (c) in the main text  [^1] When considering samples from the uniform distribution over electronic configurations on the correct particle sector, the configuration recovery method cannot be applied since it only recovers configurations with the wrong particle number.. The results for these last two sets of tests are shown in the coming subsections for the 40-, 52- and 72-qubit experiments.
V.5.1 N2 experiments
@@
Fig. S18 shows the comparison between the HPC quantum estimator run with and without configuration recovery. The ground-state energy is shown as a function of the bond length. The estimator run on $\mathcal{X}_{N}$, i.e. without configuration recovery, obtains energies comparable to RHF energies. A level of accuracy comparable to RHF is not desirable for correlated electronic structure methods. However, since no mitigation techniques have been used in the large circuits producing the samples, this is an indication that a useful quantum signal (mostly of mean-field character) is present in $\tilde{\mathcal{X}}$. As expected, the estimator with configuration recovery shows higher levels of accuracy as compared to not using configuration recoivery, obtaining a qualitatively correct dissociation curve upon increasing $d$.
The accuracy of the energies obtained by the HPC quantum estimator without configuration recovery could be improved by obtaining more measurement outcomes and thus more configurations on the right particle sector. The typical values of $d$ used for the estimator with configuration recovery are larger than $d^{hw}_{N}$. With current noise rates, obtaining a larger size of $|\mathcal{X}_{N}|$ to run the estimator without configuration recovery, requires the collection of many more measurement outcomes, at an efficiency where approximately only 1 in 1000 measured configurations has the right particle number (see Tab 2 for the precise fraction). Configuration recovery allows improving the accuracy of the HPC quantum estimator without the collection of more measurement outcomes from the quantum device. This is desirable as it reduces the overall cost of the calculation.
Figure S18: Comparison of the performance of the HPC quantum estimator, with and without configuration recovery, run on the same set of measurement outcomes from the quantum device. The dissociation of N2 (cc-pVDZ) is considered. We compare the performance of the estimator with no configuration recovery (crosses) to the performance of the estimator with configuration recovery (points). Different numbers of configurations $d$ are considered, labeled by different colors. $d_{N}^{hw}$ refers to the number of configurations on the right particle sector of $N$ electrons extracted from the measurement outcomes, used for the estimator without configuration recovery (see the text for more details). We note that with current noise rates, $d_{N}^{hw}<1.44\textrm{ M}$. Different classical methods are shown for reference, as indicated in the legend.
@@
We first compare the performance of the estimator with configuration recovery applied to measurement outcomes from the quantum device versus its application to configurations sampled from the uniform distribution in the right particle sector. The estimator is applied to $10^{5}$ samples in both scenarios and we use the same value of $d$. In both cases, we take $K=10$ batches of samples and report $\min_{k}\left(E^{(k)}\right)$ for the potential energy surface. When running the estimator with configuration recovery we use 10 self-consistent iterations. Recall that the configuration recovery procedure recovers configurations that have the wrong particle number due to noise, thus the application of the estimator to configurations sampled from the uniform distribution in the right particle sector is equivalent to applying the SCI eigenstate solver to the random configurations. Panel (a) in Fig. S19 shows the comparison. We observe that the potential energy surface obtained from the measurement outcomes from the quantum device is both qualitatively and quantitatively more accurate than the one obtained from the uniform distribution over configurations in the right particle sector. The potential energy surface obtained from the configurations sampled from the uniform distribution shows a very prominent non-smooth behavior with energies worse than the RHF energies for a number of bond lengths.
@@
In spite of this, Fig. S19 (b) shows a higher level of accuracy obtained by the estimator with configuration recovery run using measurement outcomes from the quantum device. These results indicate that the quantum device has a signal that is amplified by the configuration recovery procedure and used by the HPC quantum estimator to procedure more accurate ground-state representations than running the same estimator on uniform noise.
Figure S19: Comparison of the estimator with configuration recovery applied to measurement outcomes from the quantum device and configurations drawn from the uniform distribution, in the dissociation of N2 (cc-pVDZ). Pentagons correspond to the energies obtained with the configurations sampled from the uniform distribution while dots correspond to the energies obtained from configurations sampled from a quantum device. Different classical methods are shown for reference, as indicated in the legend. (a) Uniform samples of configurations in the right particle sector. (b) Uniform samples of configurations over Fock space.
V.5.2 [2Fe-2S] experiments
@@
Fig. S20 shows the Kernel Density Estimation (KDE) of the $E^{(k)}$ distribution running the HPC quantum estimator both with and without configuration recovery for different values of $d$. The qualitative behavior of the KDE for the estimator with and without vonfiguration recovery is similar. Upon increasing the value of $d$, the distributions shift towards lower values of the energy. However, the distributions of $E^{(k)}$ for the estimator with configuraiton recovery have lower energy than those without. Without configuration recovery, the best ground-state energy obtained is between the RHF and CISD energies, whereas when using the configuration recovery procedure, the energy is decreased even below the CCSD estimate. Again, note that the fact that the estimator without configuration recovery still has energies below RHF is a promising indication that the raw data from the device contains a signal that the configuration recovery procedure can amplify.
Figure S20: Comparison of the performance of the HPC quantum estimator, with and without configuration recovery, run on the same set of measurement outcomes for the ground state of [2Fe-2S]. The kernel density estimation (KDE) of the distribution of ground state energies obtained from the different batches of configurations is shown. The performance of the estimator with no configuration recovery (dashed lines) is compared to the performance of the estimator with configuration recovery (solid lines). Different panels correspond to different subspace dimensions $d$, as indicated in each panel. In each panel, the solid and dashed vertical lines matching the color of the KDE shows the lowest value of the energy amongst the different batches of configurations. The grey vertical lines show the value of the ground-state energy obtained from different classical methods, as indicated at the top of the panel.
@@
Figure S21: Energy-variance analysis in [2Fe-2S] of the approximate eigenstates obtained by running the estimator with configuration recovery on measurement outcomes from the quantum device (panel (a)) and from the uniform distribution over the set of configurations that have the right particle number (panel (b)). Crosses show the energy-variance relation of HCI for reference, alongside the fit to a dashed line. Dots and pentagons correspond to energy-variance points obtained from the estimator applied to different batches of configurations. The number of determinants in each batch is encoded in the color bar. The inset is a zoom into the HCI extrapolation of the second Eigenstate.
@@
These results indicate that the quantum device provides a useful signal that is used by the HPC quantum estimator to procedure more accurate ground-state representations than using the same estimator on samples from uniform distribution.

### v3

@@ 4.3 Effect of orbital optimizations in the dissociation of N2 (cc-pVDZ)
Figure S13: Effect of orbital optimizations (OO) in the performance of SQD with configuration recovery in the dissociation of N2 (cc-pVDZ). The ground-state energy is shown as a function of the bondlength. The estimator is run on measurement outcomes from a quantum processor. The energies shown correspond to the lowest energy amongst the $K$ batches of configurations, i.e. ${\min_{k}\left(E^{(k)}\right)}$. Energies from different classical methods are shown for reference, as indicated in the legend. Results are shown for the estimator run with different numbers of configurations $d$, as indicated in each panel (a)-(d).
Figure S14: Orbital occupancy profile ${\bf{n}}$ for the [2Fe-2S], produced by SQD with configuration recovery. The increasing index $p$ labels molecular orbitals of increasing energy. The occupancy profile is shown for the three eigenstates identified in the energy-variance analysis (see Fig. 4 (b) in the main text).
Figure S15: Orbital occupancy profile $\tilde{{\bf{n}}}$ for the [2Fe-2S], produced by SQD with configuration recovery. The increasing index $t$ labels localized orbitals as shown in the bottom of the panel. The occupancy profile is shown for the three eigenstates identified in the energy-variance analysis (see Fig. 4 (b) in the main text). The representation of the orbitals is obtained from the repository  (?) from Ref.  (?).
4.4 Orbital occupation numbers for the low-energy spectrum of [2Fe-2S]
The calculations reported in Figure 4 of the main text, for the [2Fe-2S] cluster, identified three distinct eigenstates that we labeled A, B, and C. In this section, we provide additional information about the nature of the three identified eigenstates. The purpose of this analysis is to illustrate the robust classification of the wavefunctions produced by our estimator into three families (A, B, and C) but also to shed light on the nature of the corresponding eigenstates and the ability of our method to accurately capture a multireference character in electronic eigenstates, pinpointing areas of possible future improvement. In particular, we show the spatial orbital occupancy for the three eigenstates that SQD finds for [2Fe-2S] both in the MO basis and in a basis of spatially local orbitals. The MO basis helps highlight deviations from the mean-field states or lack thereof, and the basis of localized orbitals (identifiable with e.g. Fe $3d$ and S $2p$ orbitals) helps characterize our method’s ability to capture local antiferromagnetic correlations between transition-metal atoms.
@@ 4.4 Orbital occupation numbers for the low-energy spectrum of [2Fe-2S]
To further understand the nature of Eigenstates B and C, we study the occupancy in the basis of localized orbitals $\tilde{{\bf{n}}}$ (see Sec. Molecular species and active spaces for details about the basis of localized orbitals). The reported occupancies are constructed as follows: first, we obtain the one-body density matrix for the approximate ground state obtained from each batch of configurations, expressed in the basis of MOs: $\Gamma_{pq,\sigma}^{(k)}$ (see Eq. S41). We then rotate the density matrix to the basis of localized orbitals with a similarity transformation denoted by $\Omega$,
@@
Fig. S15 shows a comparison of the value of $\tilde{{\bf{n}}}$ for Eigenstates A, B, and C. The reflection symmetry along the plane that separates the left and right sides of the molecule is broken for the three approximate eigenstates. This is highlighted by the observation that the occupancy of the equivalent orbitals for the two Fe atoms are not the same. We also observe that the occupancy of the bonding orbitals of the SCH3 groups and the occupancy of the $3p$ orbitals of the S atoms is similar for the three eigenstates. The occupancy of the $3d$ orbitals of Fe has the largest differences among the three different eigenstates. The more notable discrepancy in the d-orbitals of Fe may be a consequence of the fact that the low-energy physics of this molecule is dominated by the antiferromagnetic coupling of the electrons in the Fe $3d$ orbitals  (?, ?).
4.5 Quantum signal in the experiments
@@
one should verify whether there is a useful signal coming out of the quantum circuits, comparing samples from $\tilde{P}_{\Psi}$ and uniform random samples. We first investigate the fraction of sampled configurations that live in the correct particle sector $p_{s}^{\textrm{hw}}$ and compare it to the fraction that would be obtained if the samples came from the uniform distribution $p_{s}^{\textrm{unif}}$. The values for $p_{s}^{\textrm{hw}}$ alongside their 95% confidence interval, and $p_{s}^{\textrm{unif}}$ are reported in Table S1. For the 40-, 52-, and 72-qubit experiments, we observe that $p_{s}^{\textrm{hw}}$ is $\sim 20$, $\sim 10^{3}$, and $\sim 40$ times larger than $p_{s}^{\textrm{unif}}$ respectively. This provides a first indication that configurations sampled from the quantum processor contain a signal that is distinguishable from white noise.
The second set of tests that we conduct is the comparison of the ground-state energy obtained by SQD with and without applying the configuration recovery procedure. Note that not applying configuration recovery makes the estimator equivalent to the QSCI framework  (?) . Having reasonable energies (lower than or in the vicinity of the RHF energy) without applying configuration recovery reveals that the set $\mathcal{X}_{N}$ already contains reasonable configurations to construct a non-random ground state.
The third set of tests involves comparing the performance of SQD with configuration recovery run on configurations obtained from the quantum processor and samples obtained from the uniform distribution. We consider the uniform distribution over all possible configurations in the Fock space, as well as the uniform distribution over configurations with the correct particle number, as in Fig. 4 (c) in the main text  [^1] When considering samples from the uniform distribution over electronic configurations on the correct particle sector, the configuration recovery method cannot be applied since it only recovers configurations with the wrong particle number.. The results for these last two sets of tests are shown in the coming subsections for the 40-, 52- and 72-qubit experiments.
4.5.1 N2 experiments
@@
Fig. S16 shows the comparison between SQD run with and without configuration recovery. The ground-state energy is shown as a function of the bond length. The estimator run on $\mathcal{X}_{N}$, i.e. without configuration recovery, obtains energies comparable to RHF energies. A level of accuracy comparable to RHF is not desirable for correlated electronic structure methods. However, since no mitigation techniques have been used in the large circuits producing the samples, this is an indication that a useful quantum signal (mostly of mean-field character) is present in $\tilde{\mathcal{X}}$. As expected, the estimator with configuration recovery shows higher levels of accuracy as compared to not using configuration recoivery, obtaining a qualitatively correct dissociation curve upon increasing $d$.
The accuracy of the energies obtained by SQD without configuration recovery could be improved by obtaining more measurement outcomes and thus more configurations on the right particle sector. The typical values of $d$ used for the estimator with configuration recovery are larger than $d^{hw}_{N}$. With current noise rates, obtaining a larger size of $|\mathcal{X}_{N}|$ to run the estimator without configuration recovery, requires the collection of many more measurement outcomes, at an efficiency where approximately only 1 in 1000 measured configurations has the right particle number (see Tab S1 for the precise fraction). Configuration recovery allows improving the accuracy of SQD without the collection of more measurement outcomes from the quantum processor. This is desirable as it reduces the overall cost of the calculation.
Figure S16: Comparison of the performance of SQD, with and without configuration recovery, run on the same set of measurement outcomes from the quantum processor. The dissociation of N2 (cc-pVDZ) is considered. We compare the performance of the estimator with no configuration recovery (crosses) to the performance of the estimator with configuration recovery (points). Different numbers of configurations $d$ are considered, labeled by different colors. $d_{N}^{hw}$ refers to the number of configurations on the right particle sector of $N$ electrons extracted from the measurement outcomes, used for the estimator without configuration recovery (see the text for more details). We note that with current noise rates, $d_{N}^{hw}<1.44\textrm{ M}$. Different classical methods are shown for reference, as indicated in the legend.
@@
We first compare the performance of the estimator with configuration recovery applied to measurement outcomes from the quantum processor versus its application to configurations sampled from the uniform distribution in the right particle sector. The estimator is applied to $10^{5}$ samples in both scenarios and we use the same value of $d$. In both cases, we take $K=10$ batches of samples and report $\min_{k}\left(E^{(k)}\right)$ for the potential energy surface. When running the estimator with configuration recovery we use 10 self-consistent iterations. Recall that the configuration recovery procedure recovers configurations that have the wrong particle number due to noise, thus the application of the estimator to configurations sampled from the uniform distribution in the right particle sector is equivalent to applying the SCI eigenstate solver to the random configurations. Panel (a) in Fig. S17 shows the comparison. We observe that the potential energy surface obtained from the measurement outcomes from the quantum processor is both qualitatively and quantitatively more accurate than the one obtained from the uniform distribution over configurations in the right particle sector. The potential energy surface obtained from the configurations sampled from the uniform distribution shows a very prominent non-smooth behavior with energies worse than the RHF energies for a number of bond lengths.
@@
In spite of this, Fig. S17 (b) shows a higher level of accuracy obtained by the estimator with configuration recovery run using measurement outcomes from the quantum processor. These results indicate that the quantum processor has a signal that is amplified by the configuration recovery procedure and used by SQD to procedure more accurate ground-state representations than running the same estimator on uniform noise.
Figure S17: Comparison of the estimator with configuration recovery applied to measurement outcomes from the quantum processor and configurations drawn from the uniform distribution, in the dissociation of N2 (cc-pVDZ). Pentagons correspond to the energies obtained with the configurations sampled from the uniform distribution while dots correspond to the energies obtained from configurations sampled from a quantum processor. Different classical methods are shown for reference, as indicated in the legend. (a) Uniform samples of configurations in the right particle sector. (b) Uniform samples of configurations over Fock space.
4.5.2 [2Fe-2S] experiments
@@
Fig. S18 shows the Kernel Density Estimation (KDE) of the $E^{(k)}$ distribution running SQD both with and without configuration recovery for different values of $d$. The qualitative behavior of the KDE for the estimator with and without vonfiguration recovery is similar. Upon increasing the value of $d$, the distributions shift towards lower values of the energy. However, the distributions of $E^{(k)}$ for the estimator with configuraiton recovery have lower energy than those without. Without configuration recovery, the best ground-state energy obtained is between the RHF and CISD energies, whereas when using the configuration recovery procedure, the energy is decreased even below the CCSD estimate. Again, note that the fact that the estimator without configuration recovery still has energies below RHF is a promising indication that the raw data from the processor contains a signal that the configuration recovery procedure can amplify.
Figure S18: Comparison of the performance of SQD, with and without configuration recovery, run on the same set of measurement outcomes for the ground state of [2Fe-2S]. The kernel density estimation (KDE) of the distribution of ground state energies obtained from the different batches of configurations is shown. The performance of the estimator with no configuration recovery (dashed lines) is compared to the performance of the estimator with configuration recovery (solid lines). Different panels correspond to different subspace dimensions $d$, as indicated in each panel. In each panel, the solid and dashed vertical lines matching the color of the KDE shows the lowest value of the energy amongst the different batches of configurations. The grey vertical lines show the value of the ground-state energy obtained from different classical methods, as indicated at the top of the panel.
@@
Figure S19: Energy-variance analysis in [2Fe-2S] of the approximate eigenstates obtained by running the estimator with configuration recovery on measurement outcomes from the quantum processor (panel (a)) and from the uniform distribution over the set of configurations that have the right particle number (panel (b)). Crosses show the energy-variance relation of HCI for reference, alongside the fit to a dashed line. Dots and pentagons correspond to energy-variance points obtained from the estimator applied to different batches of configurations. The number of determinants in each batch is encoded in the color bar. The inset is a zoom into the HCI extrapolation of the second Eigenstate.
@@
These results indicate that the quantum processor provides a useful signal that is used by SQD to procedure more accurate ground-state representations than using the same estimator on samples from uniform distribution.